# ORBIT: Oak Ridge Base Foundation Model for Earth System Predictability

### v1

@@ Abstract
Earth system predictability is challenged by the complexity of environmental dynamics and the multitude of variables involved. Current AI foundation models, although advanced by leveraging large and heterogeneous data, are often constrained by their size and data integration, limiting their effectiveness in addressing the full range of Earth system prediction challenges. To overcome these limitations, we introduce the Oak Ridge Base Foundation Model for Earth System Predictability (ORBIT), an advanced vision-transformer model that scales up to 113 billion parameters using a novel hybrid tensor-data orthogonal parallelism technique. As the largest model of its kind, ORBIT surpasses the current climate AI foundation model size by a thousandfold. Performance scaling tests conducted on the Frontier supercomputer have demonstrated that ORBIT achieves 230 to 707 PFLOPS, with scaling efficiency maintained at 78% to 96% across 24,576 AMD GPUs. These breakthroughs establish new advances in AI-driven climate modeling and demonstrate promise to significantly improve the Earth system predictability.
@@ I Problem Overview
In response to the above challenges, AI-based models are emerging as a promising approach and aim to integrate prediction at different scales into a single system. These models leverage advanced machine learning techniques to handle the intricacies of the Earth system data, potentially revolutionizing our ability to forecast and respond to climate change. Notable models like FourCastNet [1], GraphCast [2], and Pangu-Weather [3] primarily use consistent datasets such as ERA5 [4] for weather forecasting, which limits their flexibility and applicability across varied tasks. Their reliance on homogeneous training and testing data is a significant challenge, restricting their utility for diverse Earth system predictability applications. Moreover, Earth system modeling benefits from abundant simulation data from multi-model resemble projects such as CMIP6 [5], which provide deep insights into the system’s structure, function, and dynamics. However, these rich datasets are underutilized in advancing Earth system modeling capabilities. Addressing this gap by integrating such extensive data can markedly improve AI model robustness and applicability, pushing the boundaries of what these tools can achieve in Earth sciences.
Recently, an AI foundation model based on the vision transformer (ViT) architecture, called ClimaX [6], has been proposed for both weather and climate modeling. It enables pre-training on CMIP6 multi-model simulation data and can be fine-tuned on limited labeled data for a variety of prediction tasks, including weather forecasting, sub-seasonal to seasonal prediction, and multi-decadal climate projection. Despite that ClimaX shows promising prediction performance, it has a relatively small model size (115 million parameters) and considers a small number of variables from limited atmospheric layers, which limits its capacity for accurate prediction and scalability potential on large leadership-class HPC systems.
@@
We develop ORBIT, Oak Ridge Base Foundation Model for Earth System Predictability. We designed ORBIT to (1) enable scaling the model size up to 113 billion parameters, (2) incorporate 91 channels of climate variables, and (3) pre-train on 10 different CMIP6 datasets [5] with 1.2 million observation data points. Our reported 113-billion model represents the largest dense ViTs to date, surpassing the existing largest dense vision model [7] by five-fold, and a thousand-fold larger than the largest AI foundation model for climate—ClimaX [6]. ORBIT achieves 707/230 PFLOPS for the 10/113 billion parameter models on 24,576 Frontier AMD GPUs using mixed-single BFLOAT16 precision.
@@
To scale ORBIT, we propose a novel Hybrid Sharded Tensor-Data Orthogonal Parallelism (Hybrid-STOP) by taking advantage of unique mathematical property for matrix chain multiplication and distributing model parameters among GPUs in alternating row and column shards. Through this innovation, the Hybrid-STOP algorithm combines tensor parallelism and fully sharded data parallelism together to achieve better scalability that a single parallelism cannot achieve. Importantly, this approach is designed to be architecture-agnostic, not requiring specialized hardware, which broadens its applicability across various computing platforms.
Our work represents the first instance of pre-training a large ViT model on a non-NVIDIA platform, achieving impressive scaling efficiency. This accomplishment is particularly significant given the limited interconnect capabilities of the Frontier platform, especially when compared to the faster interconnects of more recent NVIDIA platforms, and the less advanced software stack typical of non-NVIDIA environments. The successful scaling of the ORBIT model under these conditions not only highlights its robust design but also underscores the potential of alternative hardware platforms for training large AI models, thanks to the innovative Hybrid-STOP scaling approach that we developed.
@@ II State of The Art
Unfortunately, ClimaX only has 115 million model parameters. The small model size, in turn, limits ViT’s capability to learn complex data patterns from diverse and chaotic earth system information. In fact, the limitation of small vision models is not unique to climate modeling. The largest foundational dense ViT model to date has 22 billion parameters [7], and the size trails far behind those of the NLP models, which are now demonstrating capabilities in the trillion parameters scale [8, 9, 10].
@@ III-A Hybrid Sharded Tensor-Data Parallelism
The above equations are directly relevant to the core operations of the ViT training block. Each layer of the transformer training block, as shown in Fig. 1, consists of a self-attention and a feed-forward sub-layer. For both sub-layers, the primary computations can be summarized as performing two matrix multiplications in sequence. For the self-attention sub-layer, the primary computations are $\text{softmax}\left(QK^{T}\right)V$, where $Q$ is a query matrix, $K$ is a key matrix and $V$ is a value matrix [19]. For the feed-forward sub-layer, the primary computations are $\text{GeLU}\left(xA\right)B$, where $x$ is input to the feed-forward sub-layer. $A$ and $B$ are model parameters for linear transformation [19, 13]. Notice that even if we include the additional computations for softmax and GeLU, the key operations for both the self-attention and the feed-forward sub-layers are still conformed to the general form: $y\leftarrow xAB$. The Hybrid-STOP algorithm recognizes this pattern between the feed-forward and self-attention sub-layers. It uses Eqns. (2) and (3) to distribute model parameters and computations for both self-attention and feed-forward, while also integrating tensor and FSDP parallelisms.
@@
At time steps $T_{2}$ and $T_{3}$ in Fig. 3(a), the first FSDP group, consisting of GPUs 1 and 2, gathers the left half columns of the model parameters $A$ from Eqns. (2) and 3). Then the activations corresponding to the left half columns are computed at time step $T_{4}$. Similarly, the second FSDP group, consisting of GPUs 3 and 4, gathers the right half column shard of the model parameters $A$ and computes half activations. Then at time step $T_{6}$, the first FSDP group gathers the model shards for the top half row shard for the model parameters $B$ from Eqns. (2) and 3), and the second FSDP group gathers the bottom half row shard for $B$. By multiplying the half columns activation from time step $T_{4}$ with the half row shard for $B$ from time step $T_{6}$, each GPU computes an individual activation output, $xA_{*,k}B_{k,*}$ at time step $T_{7}$. Then GPUs in each tensor-parallel group add their individual activations together, based on the mathematical principle defined in Eq. (2), and compute the final activation and training loss.
In the backward pass in Fig. 3(b), each FSDP group gathers their row shards for model parameters $B$ at time step $T_{1}$. Then each FSDP group performs a gather and a reduce-scatter operations to compute the gradients for the parameter row shards at time step $T_{2}$, in the same manner as the FSDP algorithm operates in Fig. 2(b). The only difference is that hybrid-STOP does not gather the full model. Instead, GPUs 1 and 2 gathers the first half row shard, and GPUs 3 and 4 gathers the second half row shard. Then, each FSDP group gathers column shards of model parameters $A$ at step $T_{3}$ and compute the gradient for its column shard at step $T_{4}$. At time step $T_{5}$, the gradients for each GPU’s activation are computed. To synchronize gradients, GPUs in the same tensor-parallel group add their activation gradients together through all-reduce to compute the final gradients, following the mathematical principle in Eq. (3).
@@ III-B Optimization
Architecture Optimization. Prior work on training 22-billion ViT model reports divergent training loss due to extremely large attention logits with near-zero entropy [7]. To solve this issue, we adopted the same approach from [7], by applying additional layer normalization to the queries, $Q$ and keys, $V$, before computing the self-attention scaled dot product. Through these additional layer normalization, the attention logit growth is contained and the training loss divergence is prevented.
@@ IV How Performance Was Measured
Model Configuration. All scalability performance numbers presented in the Result Sec. V are based on five different ViT model sizes: 115 million (1024 embedding, 8 layers, 8 attention heads), 1 billion (3072 embedding, 8 layers, 8 attention heads), 10 billion (8192 embedding size, 11 layers, 32 attention heads) and 113 billion parameters (12288 embedding size, 56 layers, 32 attention heads). The architecture was the same as the ClimaX AI foundational model [6], with a figure illustration in Fig. 1. The only exception is that we used additional layer normalization to self-attention queries and key vectors, as described before in Sec. III-B, to prevent training loss divergence.
@@
Walltime (secs) OOM 1.83 0.49 0.47 0.33
Each pre-training observation data point has size $128\times 256\times C_{d}$, where 128×256128$\times$ 256128 × 256 is an image for a single climate variable with a spatial resolution of 1.40625 degree. $C_{d}$ is the configurable input channel dimension and represents the number of climate variables to be used for training. At each model size (115 million, 1 billion, 10 billion, and 113 billion), we pre-trained two different ViT models, one with an input channel of 48 climate variables ($C_{d}=48$) and one with an input channel of 91 variables ($C_{d}=91$). The 48 climate variables are chosen based on the variable settings in ClimaX [6]. The 91 variables are carefully chosen by climate scientists to represent an even broader range of physical variables, including 3 static variables, 3 surface variables, and 85 atmospheric variables spanning 17 pressure levels in the atmosphere. The full list of the 91 chosen variables will be provided in the final paper’s appendix.
@@
Fine-Tuning Dataset. We utilized the ERA5 dataset [4], spanning from 1979 to 2020, for model fine-tuning. Specifically, data from 1979 to 2018 was allocated for training the fine-tuning data. Data from 2019 and 2020 was used for validation and evaluation, respectively. This partition of the fine-tuning dataset aligns with the current standard used by the Weatherbench2 [25]. The fine-tuning dataset has the same spatial resolution as pre-training and uses 91 variables, with an input observation data point size of 128×256×91128$\times$ 256$\times$ 91128 × 256 × 91. The output variables selected were geopotential at 500 hPa (z500), temperature at 850 hPa (t850), 2-meter temperature (t2m), and zonal wind at 10 meters (u10).
@@
  • Time-to-solutions. We reported two time-to-solutions for pre-training, one for 48 channel variables and one for 91 channel variables. The size of each 48-channel observation data point is 128×256×48128$\times$ 256$\times$ 48128 × 256 × 48. The size of each 91-channel data point is 128×256×91128$\times$ 256$\times$ 91128 × 256 × 91. The time-to-solution is defined as the average wall clock time to process each observation data point. The time-to-solution also equals to the runtime per epoch, divided by the number of data points in an epoch.
@@
For the final model prediction performances on fine-tuning evaluation data, predictions were evaluated over the entire grid points for the year 2020. The area differences in the grid cells from the poles to the equator yield an inordinate bias to the polar regions, if all cells are weighed equally. Thus, we evaluated the prediction performance based on two latitude-weighted metrics. One metric is wRMSE, and the other metric is latitude-weighted Anomaly Correlation Coefficient (wACC). wACC is calculated as the Pearson correlation coefficient of the anomalies with respect to the climatology for each variable, with the values ranging from -1 (perfect anti-correlation) to 1 (perfect correlation). 0 represents predictions that are identical to the climatology. A higher wACC suggests that the model predictions closely follow the actual spatial variations observed in climate data, which is desirable. On the other hand, a lower wRMSE is indicative of better model accuracy.
@@ V-B Optimization Techniques
Table I shows the walltime comparison, with and without the optimizations techniques from Sec. III-B. All reported numbers in the table were for the 113 billion parameters model on 512 GPUs (64 nodes), using 48 channels of variables. The program was out of memory (OOM) when none of the optimization techniques was used. With layer wrapping alone, walltime was 1.83 seconds per observation data point. With both layer wrapping and BFLOAT16 mixed-precision, walltime was reduced to 0.49 seconds. With all four optimizations, the walltime was further reduced to 0.33 seconds.
@@ V-C Configurations for Hierarchical Parallelism
V-D Strong Scaling and Time-to-Solution
Fig. 7 shows the strong scaling efficiencies, denoted as “E” in the figure, and walltime per observation data points, denoted as “T” in the figure, from 512 GPUs (64 nodes) to 24,576 GPUs (3,072 nodes). Experiments in Fig. 7(a) used 48 channels of climate variables and experiments in Fig. 7(b) used 91 channels of climate variables. We observed that the Hybrid-STOP algorithm retained excellent strong scaling efficiency when scaling up the model size and the input channel size. With 48 input channels, all four model sizes (113 billion, 10 billion, 1 billion, and 115 million parameters) maintained strong scaling efficiencies in the range of 78% to 83% at 24,576 GPUs, compared to the baseline performance at 512 GPUs. For the performance with 91 input channels, all model sizes retained strong scaling efficiencies in the range of 81% to 96% at 24,576 GPUs.
For time-to-solution, the 113 billion parameter model takes 8E-03 second for each 48-variable observation data point at 24,576 GPUs with a sustained computing throughput at 230 PFLOPS. One pre-training epoch takes 2.7 wall-clock hours on 24,576 GPUs for the 113 billion parameter model to see all 1.2 million observation data points. The 10 billion parameter model takes 3E-04 second for each 48-channel observation data point at 24,576 GPUs, with a sustained computing throughput at 707 PFLOPS. In contrast, due to increased memory use for more input channels, processing each 91-channel observation data point takes more walltime, requiring 1E-02 second for the 113 billion parameter model, and 4E-04 second for the 10 billion parameter model at 24,576 GPUs.

### v3

@@ Abstract
Earth system predictability is challenged by the complexity of environmental dynamics and the multitude of variables involved. Current AI foundation models, although advanced by leveraging large and heterogeneous data, are often constrained by their size and data integration, limiting their effectiveness in addressing the full range of Earth system prediction challenges. To overcome these limitations, we introduce the Oak Ridge Base Foundation Model for Earth System Predictability (ORBIT), an advanced vision transformer model that scales up to 113 billion parameters using a novel hybrid tensor-data orthogonal parallelism technique. As the largest model of its kind, ORBIT surpasses the current climate AI foundation model size by a thousandfold. Performance scaling tests conducted on the Frontier supercomputer have demonstrated that ORBIT achieves 684 petaFLOPS to 1.6 exaFLOPS sustained throughput, with scaling efficiency maintained at 41% to 85% across 49,152 AMD GPUs. These breakthroughs establish new advances in AI-driven climate modeling and demonstrate promise to significantly improve the Earth system predictability.
@@ I Problem Overview
In response to the above challenges, AI-based models are emerging as a promising approach and aim to integrate prediction at different scales into a single system. These models leverage advanced machine learning techniques to handle the intricacies of the Earth system data, potentially revolutionizing our ability to forecast and respond to climate change. Notable task-specific models like FourCastNet [1], GraphCast [2], Pangu-Weather [3], and Stormer [4] primarily use consistent datasets such as ERA5 [5] for weather forecasting, which limits their flexibility and applicability across varied tasks. Their reliance on homogeneous training and testing data is a significant challenge, restricting their utility for diverse Earth system predictability applications. Moreover, Earth system modeling benefits from abundant simulation data from multi-model resemble projects such as CMIP6 [6], which provide deep insights into the system’s structure, function, and dynamics. However, these rich datasets are underutilized in advancing Earth system modeling capabilities. Addressing this gap by integrating such extensive data can markedly improve AI model robustness and applicability, pushing the boundaries of what these tools can achieve in Earth sciences.
Recently, an AI foundation model based on the vision transformer (ViT) architecture, called ClimaX [7], has been proposed for both weather and climate modeling. It enables pre-training on CMIP6 multi-model simulation data and can be fine-tuned on limited labeled data for a variety of prediction tasks, including weather forecasting, sub-seasonal to seasonal prediction, and multi-decadal climate projection. Despite of ClimaX’s promising prediction performance, it has a relatively small model size of 115 million parameters and considers a limited number of variables (also known as channels) from selected atmospheric layers. These limitations, in turn, constrain the ClimaX model’s capacity for accurate prediction and scalability potential on large leadership-class HPC systems.
@@
We develop ORBIT, Oak Ridge Base Foundation Model for Earth System Predictability. We designed ORBIT to (1) enable scaling the model size up to 113 billion parameters with throughput up to 1.6 exaFLOPS, (2) incorporate 91 channels of climate variables, and (3) pre-train on 10 different CMIP6 datasets [6] with 1.2 million observation data points. In comparison, the current state-of-the-art ClimaX foundation model uses 115 million parameters, 48 variables and 5 CMIP6 datasets. Our reported 113-billion model represents the largest dense ViTs to date, surpassing the existing largest dense vision model [8] by five-fold, and a thousand-fold larger than the largest AI foundation model for climate—ClimaX [7] and two hundred times larger than domain-specific Stormer model [4]. ORBIT achieves 1.6 exaFLOPS/684 PFLOPS for the 10/113 billion parameter models on 49,152 Frontier AMD GPUs using mixed-single BFLOAT16 precision.
@@
To scale ORBIT, we propose a novel Hybrid Sharded Tensor-Data Orthogonal Parallelism (Hybrid-STOP) by taking advantage of unique mathematical property for matrix chain multiplication and distributing model parameters among GPUs in alternating row and column shards. Through this innovation, the Hybrid-STOP algorithm combines tensor parallelism and fully sharded data parallelism together to achieve better scalability that a single parallelism cannot achieve. Importantly, this approach is designed to be architecture-agnostic, not requiring a specific hardware, which broadens its applicability across various computing platforms.
Our work represents the first instance of pre-training a large ViT model on a non-NVIDIA platform, achieving impressive scaling efficiency and exaFLOPS sustained throughput. This accomplishment is particularly significant given the limited interconnect capabilities of the Frontier platform, especially when compared to the faster interconnects of more recent NVIDIA platforms, and the less advanced software stack typical of non-NVIDIA environments. The successful scaling of the ORBIT model under these conditions not only highlights its robust design but also underscores the potential of alternative hardware platforms for training large AI models, thanks to the innovative Hybrid-STOP scaling approach that we developed.
@@ II State of The Art
ClimaX only has 115 million model parameters. The small model size, in turn, limits ViT’s capability to learn complex data patterns from diverse and chaotic earth system information. In fact, the limitation of small vision models is not unique to climate modeling. The largest foundational dense ViT model to date has 22 billion parameters [8], and the size trails far behind those of the NLP models, which are now demonstrating capabilities in the trillion parameters scale [9, 10, 11].
@@ III-A Hybrid Sharded Tensor-Data Parallelism
The above equations are directly relevant to the core operations of the ViT training block. Each layer of the transformer training block, as shown in Fig. 1, consists of a self-attention and a feed-forward sub-layer. For both sub-layers, the primary computations can be summarized as performing two matrix multiplications in sequence. For the self-attention sub-layer, the primary computations are $\text{softmax}\left(QK^{T}\right)V$, where $Q$ is a query matrix, $K$ is a key matrix and $V$ is a value matrix [20]. For the feed-forward sub-layer, the primary computations are $\text{GeLU}\left(xA\right)B$, where $x$ is input to the feed-forward sub-layer. $A$ and $B$ are model parameters for linear transformation [20, 14]. Notice that even if we include the additional computations for softmax and GeLU activation, the key operations for both the self-attention and the feed-forward sub-layers are still conformed to the general form: $y\leftarrow xAB$. The Hybrid-STOP algorithm recognizes this pattern between the feed-forward and self-attention sub-layers. It uses Eqns. (2) and (3) to distribute model parameters and computations for both self-attention and feed-forward, while also integrating tensor and FSDP parallelisms.
@@
At time steps $T_{2}$ and $T_{3}$ in Fig. 3(a), the first FSDP group, consisting of GPUs 1 and 2, gathers the left half columns of the model parameters $A$ from Eqns. (2) and (3). Then the activations corresponding to the left half columns are computed at time step $T_{4}$. Similarly, the second FSDP group, consisting of GPUs 3 and 4, gathers the right half column shard of the model parameters $A$ and computes half activations. Then at time step $T_{6}$, the first FSDP group gathers the model shards for the top half row shard for the model parameters $B$ from Eqns. (2) and (3), and the second FSDP group gathers the bottom half row shard for $B$. By multiplying the half columns activation from time step $T_{4}$ with the half row shard for $B$ from time step $T_{6}$, each GPU computes an individual activation output, $xA_{*,k}B_{k,*}$ at time step $T_{7}$. Then GPUs in each tensor-parallel group add their individual activations together, based on the mathematical principle defined in Eqn. (2), and compute the final activation and training loss.
In the backward pass in Fig. 3(b), each FSDP group gathers their row shards for model parameters $B$ at time step $T_{1}$. Then each FSDP group performs a gather and a reduce-scatter operations to compute the gradients for the parameter row shards at time step $T_{2}$, in the same manner as the FSDP algorithm operates in Fig. 2(b). The only difference is that hybrid-STOP does not gather the full model. Instead, GPUs 1 and 2 gathers the first half row shard, and GPUs 3 and 4 gathers the second half row shard. Then, each FSDP group gathers column shards of model parameters $A$ at step $T_{3}$ and compute the gradient for its column shard at step $T_{4}$. At time step $T_{5}$, the gradients for each GPU’s activation are computed. To synchronize gradients, GPUs in the same tensor-parallel group add their activation gradients together through all-reduce to compute the final gradients, following the mathematical principle in Eqn. (3).
@@ III-B Optimization
Architecture Optimization. Prior work on training 22-billion ViT model reports divergent training loss due to extremely large attention logits with near-zero entropy [8]. To solve this issue, we adopted the same approach from [8], by applying additional layer normalization to the queries, $Q$, and keys, $V$, before computing the self-attention scaled dot product. Through these additional layer normalization, the attention logit growth is contained and the training loss divergence is prevented.
@@ IV How Performance Was Measured
Model Configuration. All scalability performance numbers presented in the Result Sec. V are based on five different ViT model sizes: 115 million (1024 embedding, 8 layers, 16 attention heads), 1 billion (3072 embedding, 8 layers, 16 attention heads), 10 billion (8192 embedding size, 11 layers, 32 attention heads) and 113 billion parameters (12288 embedding size, 56 layers, 64 attention heads). The architecture was the same as the ClimaX AI foundational model [7], with a figure illustration in Fig. 1. The only exception is that we used additional layer normalization to self-attention queries and key vectors, as described before in Sec. III-B, to prevent training loss divergence.
@@
Walltime (secs) OOM 0.97 0.49 0.40 0.17
Each pre-training observation data point has size $128\times 256\times C_{d}$, where $128\times 256$ is an image for a single climate variable with a spatial resolution of 1.40625 degree. $C_{d}$ is the configurable input channel dimension and represents the number of climate variables to be used for training. At each model size (115 million, 1 billion, 10 billion, and 113 billion), we pre-trained two different ViT models, one with an input channel of 48 climate variables ($C_{d}=48$) and one with an input channel of 91 variables ($C_{d}=91$). The 48 climate variables are chosen based on the variable settings in ClimaX [7]. The 91 variables are carefully chosen by climate scientists to represent an even broader range of physical variables, including 3 static variables, 3 surface variables, and 85 atmospheric variables spanning 17 pressure levels in the atmosphere. The full list of the 91 chosen variables will be provided to readers when the models are released on GitHub.
@@
Fine-Tuning Dataset. We utilized the 1.40625-degree ERA5 dataset [5], spanning from 1979 to 2020, for model fine-tuning. Specifically, data from 1979 to 2018 was allocated for training the fine-tuning data. Data from 2019 and 2020 was used for validation and evaluation, respectively. This partition of the fine-tuning dataset aligns with the current standard used by the Weatherbench2 [26]. The fine-tuning dataset has the same spatial resolution as pre-training and uses 91 variables, with an input observation data point size of $128\times 256\times 91$. The output variables selected were geopotential at 500 hPa (z500), temperature at 850 hPa (t850), 2-meter temperature (t2m), and zonal wind at 10 meters (u10).
@@
  • Time-to-solutions. We reported two time-to-solutions for pre-training, one for 48 channel variables and one for 91 channel variables. The size of each 48-channel observation data point is $128\times 256\times 48$. The size of each 91-channel data point is $128\times 256\times 91$. The time-to-solution is defined as the average wall clock time to process each observation data point. The time-to-solution also equals to the runtime per epoch, divided by the number of data points in an epoch.
@@
For the final model prediction performances on fine-tuning evaluation data, predictions were evaluated over the entire grid points for the year 2020. The area differences in the grid cells from the poles to the equator yield an inordinate bias to the polar regions, if all cells are weighed equally. Thus, we evaluated the prediction performance based on latitude-weighted Anomaly Correlation Coefficient (wACC). wACC is calculated as the Pearson correlation coefficient of the anomalies with respect to the climatology for each variable, with the values ranging from -1 (perfect anti-correlation) to 1 (perfect correlation). Value of 0 represents predictions identical to the climatology. A higher wACC suggests that the model predictions closely follow the actual spatial variations observed in climate data, which is desirable.
@@ V-B Optimization Techniques
Table I shows the walltime comparison, with and without the optimizations techniques from Sec. III-B. All reported numbers in the table were for the 113 billion parameters model on 512 GPUs (64 nodes), using 48 channels of variables. The program was out of memory (OOM) when none of the optimization techniques was used. With layer wrapping alone, walltime was 0.97 seconds per observation data point. With both layer wrapping and BFLOAT16 mixed-precision, walltime was reduced to 0.49 seconds. With all four optimizations, the walltime was further reduced to 0.14 seconds.
@@ V-C Configurations for Hierarchical Parallelism
V-D Strong Scaling and Time-to-Solutions
Fig. 7 shows the strong scaling efficiencies, denoted as “E” in the figure, and walltime per observation data points, denoted as “T” in the figure, from 512 GPUs (64 nodes) to 49,152 GPUs (6,144 nodes). Experiments in Fig. 7(a) used 48 channels of climate variables and experiments in Fig. 7(b) used 91 channels of climate variables. We observed that the Hybrid-STOP algorithm retained excellent strong scaling efficiency when scaling up the model size and the input channel size. With 48 input channels, all four model sizes (113 billion, 10 billion, 1 billion, and 115 million parameters) maintained strong scaling efficiencies in the range of 44% to 82% at 49,152 GPUs, compared to the baseline performance at 512 GPUs. For the performance with 91 input channels, all model sizes retained strong scaling efficiencies in the range of 41% to 85% at 49,152 GPUs.
For time-to-solutions, the 113 billion parameter model takes 3E-03 second for each 48-variable observation data point at 49,152 GPUs with a sustained computing throughput at 684 PFLOPS. One pre-training epoch takes 0.8 wall-clock hours on 49,152 GPUs for the 113 billion parameter model to see all 1.2 million observation data points. The 10 billion parameter model takes 1E-04 second for each 48-channel observation data point at 49,152 GPUs, with a sustained computing throughput at 1.6 exaFLOPS. In contrast, due to increased memory use for more input channels, processing each 91-channel observation data point takes more walltime, requiring 5E-03 second for the 113 billion parameter model, and 2E-04 second for the 10 billion parameter model at 49,152 GPUs.